1. Consider the following sample spot speed data. The speeds are in miles per hour Speed Grou Lower Limit Upper Limit Mid-point Frequenc 27.5 32.5 37.5 42.5 47.5 52.5 57.5 62.5 67.5 72.5 77.5 82.5 87.5 92.5 25 30 35 40 45 50 22.6 27.6 32.6 37.6 42.6 47.6 52.6 57.6 62.6 67.6 72.6 77.6 82.6 87.6 2 14 25 60 65 70 75 80 85 90 4 (a) Calculate the sample mean speed (b) Calculate the sample median speed (c) Calculate the sample modal speed (d) Calculate the sample variance (e) If the speed limit on this highway was 70 MPH, what percentage of drivers was exceeding this speed limit? (f) Determine the 50th percentile speed Assume that the population standard deviation of speeds on this highway is known to be 13 mph, conduct the following analysis: (g) Find the 95 percent two-sided confidence interval on the mean speed (h) Find the 90 percent lower-sided confidence interval on the mean speed (i) Find the 90 percent upper-sided confidence interval on the mean speed (j) IfPCX 57 mph)-?. What is ??
Answer #1

a) Sample mean = 8267 58.22

x = midpoint

f= frequency

b) \frac{N}{2}=\frac{142}{2}=121

2060 3 3 4 2 2 2 4 1 uu 3 4 ue 3 | | | 261122|1|1 4851-84-8642 5-0-5-0-5-0-5 2-3-3-4-4-5-5 6-7-7-8-8-9 し55555555555555 r772 |

(67.6,72.5) is the median group

(n/2) -B X w Estimated Median = L+ where: .L is the lower class boundary of the group containing the median . n is the totalmedian= 67.6+\frac{121-108}{14}(4.9)=72.15

c)fm-m-1 Estimated Mode = L+ m-1 where: the modal group (the group with the highest frequency) .L is the lower dass boundary of

(52.6,57.5) is the modal group

mode=52.6+\frac{25-18}{(25-18)+(25-21)}(4.9)=52.6+3.12=55.72

d) S^2=\frac{1}{n-1}[\sum fX^2-\frac{(\sum fX)^2}{n}]=\frac{1}{142-1}[512550-\frac{(8267)^2}{142}]=221.70